# Cleaning the Pool: Progressive Filtering of Unlabeled Pools in Deep Active Learning

### v1

@@ 1 Introduction
Part of the challenge is that different selection strategies capture fundamentally different notions of data value, e.g., uncertainty-based strategies target difficult instances while representativeness-based strategies seek high-density ones. Each perspective offers reasonable insights, yet no single strategy dominates across the entire AL process, as instances’ impact on model performance strongly varies per cycle [11, 4]. As the model learns, the optimal selection criterion shifts: a strategy that excels in one AL cycle may become suboptimal in the next. This means that committing to any single strategy requires accepting its blind spots. While finding a universally optimal strategy is challenging, we presume that combining multiple complementary strategies in an ensemble AL method provides a more robust alternative. While leveraging collective strengths, we reduce the risk of committing to any potentially suboptimal heuristic.
@@
Specifically, in each AL cycle, Refine operates in two stages (Fig. 1): (1) Progressive filtering iteratively refines the unlabeled pool through multiple rounds. In each round, we apply all strategies from the ensemble on the current unlabeled pool and take the union of the selected instances, retaining batches that at least one strategy considers valuable. Repeating this process yields an high-value candidate pool, i.e., it reflects complementary notions of data value while filtering out uninformative instances that provide no value according to any strategy. This pool cleaning process is a powerful preprocessing step on its own, already considerably improving the performance of individual strategies. (2) Coverage-based selection then selects the final batch from the refined pool by choosing the instances that, when combined with the already labeled pool, maximize coverage over the refined pool. Since progressive filtering has already identified high-value candidates, this selection ensures the batch is accounting for the various notions of data value captured in the filtering stage.
@@
  • We introduce progressive filtering, an approach to iteratively refine the unlabeled pool by retaining valuable instances while cleaning out uninformative ones, thereby reducing pool size and improving pool quality for downstream selection.
@@ 2 Related Work
Ensemble AL methods use multiple AL strategies by either combining them or choosing when to use which. Although ALBL [14], COMB [5], and DUAL [8] fit this regime, we do not consider them as they are single instance strategies whose adaptation to the batch setting is not straightforward. SelectAL [11] chooses between low- and high-budget strategies by evaluating which one identifies the most valuable batch of the current labeled pool. However, consistent with Bae et al. [4], we found it yields inconsistent performance and is difficult to reproduce. TAILOR [31] frames the choice of AL strategies as a multi-armed bandit problem, learning a probability distribution used to sample which strategies propose instances. As TAILOR uses a class-balance-based reward to update this distribution, it can handle class-imbalanced settings effectively. TCM [9] initially applies TypiClust for representativeness before transitioning to Margin for uncertainty, determining the switch point via a budget-dependent heuristic. This design is based on empirical findings that representative and diverse instances perform well on low budgets, while uncertainty is well-suited for high budgets [11].
@@ 4.3 Theoretical Analysis of Progressive Filtering
Theorem 3 guarantees that the pool’s average value is monotonically non-decreasing, i.e., it is guaranteed to improve or stay the same with each round. The key assumption, that AL strategies favor higher-value instances, is natural for any reasonable selection heuristic. This theorem provides the theoretical view for why progressive filtering concentrates the pool toward high-value instances. In each round, lower-value instances are filtered out at higher rates than higher-value ones, thereby shifting the distribution of the pool toward instances with greater true value. Together with the other theorems, this ensures efficient convergence to a high-quality candidate pool for the second stage.
@@ 5 Experimental Evaluation
We evaluate Refine on image classification tasks, focusing primarily on standard benchmarks (e.g., CIFAR-10). We also present a use case study on audio spectrogram classification in Appendix B. We start by establishing the experimental setup, then demonstrate that Refine achieves superior performance against state-of-the-art strategies, followed by ablation studies analyzing key design choices. Our implementation can be found at anonymous. [^1] Our code will be made publicly available upon acceptance.
@@ 5.1 Experimental Setup
Table 1: Overview of datasets, showing number of instances, number of classes $K$ and batch size $b$. CIFAR-10* was used for hyperparameter selection.
@@
Baseline Strategies. We compare Refine against random sampling and 8 state-of-the-art AL strategies. Our baselines include: (i) the uncertainty-based strategies Margin [29] and BADGE [2], (ii) the diversity-based strategies TypiClust [12] and MaxHerding [3], and (iii) the hybrid approaches BAIT [1, 15], ALFAMix [23], DropQuery [10] and UHerding [4]. Finally, we include SelectAL [11], TCM [9], and TAILOR [31], recent ensemble AL methods that combine multiple AL strategies.
Evaluation Metrics. We assess AL performance using two metrics: (i) relative learning curves, showing accuracy gains (absolute percentage points) over random sampling (from ${\mathcal{U}}_{t}$) at each cycle, and (ii) area under the learning curve (AULC), measuring cumulative performance across all cycles. Absolute learning curves are provided in Appendix D. All results are averaged over $10$ independent trials and standard errors are omitted for visibility. Additionally, we report pairwise win rates between strategies: the percentage of trials where strategy $i$ achieves higher AULC than strategy $j$ across all cycles.
@@ 5.2 Main Results
Additionally, Fig. 3 shows an exemplary subset of relative learning curves on the four most complex datasets, focusing on DINOv3 as the newest model and CLIP for its differing training process (i.e., student-teacher vs. contrastive learning). The complete set of learning curves (relative and absolute) can be found in Appendix D. We observe that Refine consistently outperforms all strategies on CIFAR-100, Food101, and Tiny ImageNet across all backbones, demonstrating its robustness across different datasets and models.
@@
Performance against Ensemble Methods. Figure 4 extends the pairwise comparison to ensemble AL methods. Here, we see that Refine clearly outperforms all other methods, achieving a 100% win rate against both SelectAL and TAILOR, and 98% against TCM. These results indicate that Refine consistently combines AL strategies more effectively than competitors. The complete set of associated learning curves can be found in Appendix D.
@@ 6 Conclusion
We introduced Refine, an ensemble AL method that effectively combines different notions of value (e.g., uncertainty and representativeness) to achieve state-of-the-art performance. Our core contribution, progressive filtering, iteratively refines the unlabeled pool via the union of batches from multiple AL strategies. This process cleans the unlabeled pool by discarding uninformative instances, yielding a high-value candidate pool for downstream selection. A subsequent coverage-based selection then effectively yields a diverse batch from this refined, high-quality pool.
Limitations. The effectiveness of Refine depends on the quality and composition of its strategy ensemble. If these strategies are highly correlated (e.g., all share a focus on diversity), the filtered pool will inherit this bias, diminishing the benefits of the ensemble. Moreover, the union can be sensitive if there are many poorly-designed strategies. As these will repeatedly select uninformative instances, this may slightly degrade the value of the candidate pool. We therefore recommend building $\mathcal{S}$} from a small set of well-established, complementary AL strategies rather than arbitrarily including numerous ones.
@@ Appendix B Use Case: Audio Spectrogram Classification
Appendix C Ablation Studies: Additional Ablations
In Fig. 5(a), we reported improvements in AULC [%], which showed that all strategies applied to the progressively filtered pool $\mathcal{C}_{R}$ lead to performance improvement. For these strategies, Fig. 9 presents the corresponding accuracy learning curves. The figure compares strategies applied to the filtered pool versus the unfiltered pool.
Figure 9 demonstrates that progressive filtering acts as a powerful preprocessing step, consistently improving the accuracies of individual AL strategies. This behavior is consistent across both DINOv2 and DINOv3 backbones. While the impact is most critical for poorly performing AL strategies, progressive filtering still enhances strong performers such as UHerding, even if the improvement is less pronounced. Notably, filtering significantly improves even random sampling, demonstrating its ability to clean uninformative instances from the unlabeled pool. Furthermore, progressive filtering can act as a crucial stabilizer for failing AL strategies. For example, while standard BAIT performs worse than random sampling on DINOv3, applying filtering fixes this failure and substantially improves the strategy’s effectiveness.
@@ Appendix C Ablation Studies: Additional Ablations
Appendix D Main Results: All Learning Curves
This section presents the complete set of learning curves used to derive Figs. 2 and 4, which were omitted from the main paper due to space limitations. For individual AL strategies, Fig. 10 displays the absolute learning curves (accuracy), while Fig. 11 shows the relative learning curves (accuracy relative to random sampling from ${\mathcal{U}}_{t}$). Similarly, for ensemble AL methods, Figs. 12 and 13 present the absolute and relative learning curves, respectively. These results provide a comprehensive overview supporting the conclusions drawn in the main text.
Figures 10 and 11 demonstrate that Refine outperforms all other individual AL strategies across nearly every dataset and backbone, including DINOv2, CLIP, and DINOv3. It excels in both low- and high-budget scenarios, consistently delivering high accuracies throughout the entire AL process. Furthermore, Refine proves to be robust, maintaining a strong lead across diverse datasets while remaining invariant to the choice of the backbone. Similarly, Figures 12 and 13 demonstrate that Refine outperforms all ensemble AL methods. These approaches generally struggle to effectively combine multiple strategies, often failing to exceed the performance of individual AL strategies. With Refine, we propose the first ensemble AL method capable of effectively combining multiple AL strategies into a unified framework that consistently surpasses individual baselines.
Figure 10: Complete set of absolute learning curves for individual AL strategies showing the accuracy for all datasets and backbone combinations.
Figure 11: Complete set of relative learning curves for individual AL strategies showing the accuracy relative to random sampling from ${\mathcal{U}}_{t}$ for all datasets and backbone combinations.
@@ Appendix D Main Results: All Learning Curves
Figure 12: Complete set of absolute learning curves for ensemble AL methods showing the accuracy for all datasets and backbone combinations.

### v2

@@ 1 Introduction
Part of the challenge is that different selection strategies capture fundamentally different notions of data value, e.g., uncertainty-based strategies target difficult instances while representativeness-based strategies seek high-density ones. Each perspective offers reasonable insights, yet no single strategy dominates across the entire AL process, as instances’ impact on model performance strongly varies per cycle [12, 4]. As the model learns, the optimal selection criterion shifts: a strategy that excels in one AL cycle may become suboptimal in the next. This means that committing to any single strategy requires accepting its blind spots. While finding a universally optimal strategy is challenging, we presume that combining multiple complementary strategies in an ensemble AL method provides a more robust alternative [21]. While leveraging collective strengths, we reduce the risk of committing to any suboptimal heuristic.
@@
Specifically, in each AL cycle, Refine operates in two stages (Fig. 1): (1) Progressive filtering iteratively refines the unlabeled pool through multiple rounds. In each round, we apply all strategies from the ensemble on the current unlabeled pool and take the union of the selected instances, retaining batches that at least one strategy considers valuable. Repeating this process yields a high-value candidate pool, i.e., it reflects complementary notions of data value while filtering out uninformative instances that provide no value according to any strategy. This pool cleaning process is a powerful preprocessing step on its own, already considerably improving the performance of individual strategies. (2) Coverage-based selection then selects the final batch from the refined pool by choosing the instances that, when combined with the already labeled pool, maximize coverage over the refined pool. Since progressive filtering has already identified high-value candidates, this selection ensures the batch is accounting for the various notions of data value captured in the filtering stage.
@@
  • We introduce progressive filtering to iteratively refine the unlabeled pool by retaining valuable instances while cleaning out uninformative ones, thereby reducing pool size and improving its quality for downstream selection.
@@ 2 Related Work
Ensemble AL methods use multiple AL strategies by either combining them or choosing when to use which. Although ALBL [17], COMB [5], and DUAL [8] fit this regime, we do not consider them as they are single instance strategies whose adaptation to the batch setting is not straightforward. SelectAL [12] chooses between low- and high-budget strategies by evaluating which one identifies the most valuable batch of the current labeled pool. However, consistent with Bae et al. [4], we found it yields inconsistent performance and is difficult to reproduce. TAILOR [40] frames the choice of AL strategies as a multi-armed bandit problem, learning a probability distribution used to sample which strategies propose instances. As TAILOR uses a class-balance-based reward to update this distribution, it can handle class-imbalanced settings effectively. TCM [9] initially applies TypiClust for representativeness before transitioning to Margin for uncertainty, determining the switch point via a budget-dependent heuristic. This design is based on empirical findings that representative and diverse instances perform well on low budgets, while uncertainty is well-suited for high budgets [12]. Finally, AutoAL [37] learns to weight the votes of multiple strategies via a model and selects the top-$k$ instances based on the total votes.
@@ 4.3 Theoretical Analysis of Progressive Filtering
Theorem 3 guarantees that the pool’s average value is monotonically non-decreasing, i.e., it is guaranteed to improve or stay the same with each round. The key assumption, that AL strategies favor higher-value instances, is natural for any reasonable selection heuristic, though it may not always hold in practice. This theorem provides the theoretical view for why progressive filtering concentrates the pool toward high-value instances. In each round, lower-value instances are filtered out at higher rates than higher-value ones, thereby shifting the distribution of the pool toward instances with greater true value. Together with the other theorems, this ensures efficient convergence to a high-quality candidate pool for the second stage.
@@ 5 Experimental Evaluation
We evaluate Refine on image classification tasks, focusing primarily on standard benchmarks (e.g., CIFAR-10). We also present a use case study on audio spectrogram classification in Appendix B. We start by establishing the experimental setup, then demonstrate that Refine achieves superior performance against state-of-the-art strategies, followed by ablation studies analyzing key design choices. Our implementation can be found at https://github.com/dhuseljic/dal-toolbox.
@@ 5.1 Experimental Setup
Table 1: Overview of datasets. CIFAR-10* was used for hyperparameter selection.
@@
Baseline Strategies. We compare Refine against random sampling and 8 state-of-the-art AL strategies. Our baselines include: (i) the uncertainty-based strategies Margin [34] and BADGE [2], (ii) the representativeness-based strategies TypiClust [13] and MaxHerding [3], and (iii) the hybrid approaches BAIT [1, 19], ALFAMix [27], DropQuery [11] and UHerding [4]. Finally, we include SelectAL [12], TCM [9], TAILOR [40], and AutoAL [37], recent ensemble AL methods that combine multiple AL strategies.
Evaluation Metrics. We assess AL performance using two metrics: (i) relative learning curves, showing accuracy gains (absolute percentage points) over random sampling (from ${\mathcal{U}}_{t}$) at each cycle, and (ii) area under the learning curve (AULC), measuring cumulative performance across all cycles. Absolute learning curves are provided in Appendix E. All results are averaged over $10$ independent trials with shaded areas representing standard errors. Additionally, we report pairwise win rates between strategies: the percentage of trials where strategy $i$ achieves higher AULC than strategy $j$ across all cycles.
@@ 5.2 Main Results
Additionally, Fig. 3 shows an exemplary subset of relative learning curves on the four most complex datasets, focusing on DINOv3 as the newest model and CLIP for its differing training process (i.e., student-teacher vs. contrastive learning). The complete set of learning curves (relative and absolute) can be found in Appendix E. We observe that Refine consistently outperforms all strategies on CIFAR-100, Food101, and Tiny ImageNet across all backbones, demonstrating its robustness across different datasets and models.
@@
Performance against Ensemble Methods. Figure 4 extends the pairwise comparison to ensemble AL methods. Refine clearly outperforms all competitors, achieving a 100% win rate against both SelectAL and TAILOR, 98% against TCM, and 97% against AutoAL, indicating that it combines AL strategies more effectively. Additional experiments and the complete set of associated learning curves can be found in Appendices D and E, respectively.
@@ 6 Conclusion
We introduced Refine, an ensemble AL method that combines different notions of value to achieve state-of-the-art performance. Our core contribution, progressive filtering, iteratively refines the unlabeled pool via the union of batches from multiple AL strategies. This process cleans the unlabeled pool by discarding uninformative instances, yielding a high-value candidate pool for downstream selection. A subsequent coverage-based selection effectively yields a diverse batch from this refined, high-quality pool.
Limitations. The effectiveness of Refine depends on the quality and composition of its strategy ensemble. If these strategies are correlated (e.g., all focus on diversity), the filtered pool will inherit this bias. Moreover, the union can be sensitive if there are many poorly-designed strategies, which may slightly degrade the value of the candidate pool. We therefore recommend building $\mathcal{S}$} from a small set of well-established, complementary strategies rather than arbitrarily including numerous ones.
@@ Appendix B Use Case: Audio Spectrogram Classification
Appendix C Additional Ablations
Filtering as Pool Refinement. In Fig. 5(a), we reported improvements in AULC [%], which showed that all strategies applied to the progressively filtered pool $\mathcal{C}_{R}$ lead to performance improvement. For these strategies, Fig. 9 presents the corresponding accuracy learning curves. The figure compares strategies applied to the filtered pool versus the unfiltered pool.
Figure 9 demonstrates that progressive filtering acts as a powerful preprocessing step, consistently improving the accuracies of individual AL strategies. This behavior is consistent across both DINOv2 and DINOv3 backbones. While the impact is most substantial for poorly performing AL strategies such as AlfaMix, progressive filtering still enhances strong performers such as UHerding, even if the improvement is less pronounced. Moreover, filtering significantly improves even random sampling, demonstrating its ability to clean uninformative instances from the unlabeled pool. Furthermore, progressive filtering can act as a crucial stabilizer for failing AL strategies. For example, while standard BAIT performs worse than random sampling on DINOv3, applying filtering fixes this issue and substantially improves the strategy’s effectiveness.
@@ Appendix C Additional Ablations
Weak Ensemble Members. We investigate a potential failure case in Refine by considering an ensemble where the majority of strategies perform worse than random sampling. Specifically, we run Refine on Dopanim (CLIP) using an ensemble of three weak and one strong strategy. All ensemble members and Refine are shown in Fig. 10 (left). Despite this unfavorable composition, Refine performs well, closely following the strong ensemble member. A promising future direction may be to measure the similarity of ensemble selections and reduce the influence of non-complementary strategies over the AL cycles. This would encourage more independent selection within the ensemble, increasing the likelihood of retaining valuable instances.
Varying Batch Size. We investigate Refine’s behavior when changing the batch size $b$. Hence, we evaluate Refine on CIFAR-100 (DINOv2) with batch sizes $b\in\{10,25,50,100\}$ under a fixed total budget of $B\approx 2500$. Interestingly, results in Fig. 10 (right) show that larger batch sizes yield slightly better performance. We attribute this to the filtering process benefiting from a larger candidate set: each strategy contributes more instances, increasing the diversity of the filtered pool. Over multiple rounds, this effectively exposes the filtering process to a larger portion of the unlabeled pool, making it more likely that valuable instances are preserved.
Figure 10: (Left) Relative learning curves of Refine and all ensemble members on Dopanim. (Right) Absolute learning curves of Refine on CIFAR-100 for different batch sizes $b$.
Alternative for second stage. While coverage-based selection in the second stage of Refine performs well empirically, it is solely a single strategy applied to the refined pool ${\mathcal{C}}_{R}$ without focusing on the most valuable heuristic in the current cycle. To provide an outlook on how we can realize this focus, we additionally consider a more deliberate selection process. Rather than applying a single coverage-based strategy to ${\mathcal{C}}_{R}$, we use each strategy in the ensemble independently. With an ensemble of two strategies, this results in two proposed batches, each reflecting a different heuristic. The selection process then reduces to assessing which of the proposed batches is expected to yield the greatest improvement in model performance.
To rank the candidate batches, we investigate two established criteria from the literature: the Fisher information ratio (FIR) [36] and the expected error (ER) [32]. Both criteria estimate the impact on model performance of including a batch in the training dataset. Notably, these criteria cannot be straightforwardly applied in a deep batch AL setting [19, 20], as the search for a candidate batch is combinatorially infeasible given a large unlabeled pool. Letting each strategy in the ensemble construct a candidate batch effectively guides the search through this otherwise intractable space [21]. The results in Fig. 11 demonstrate that this selection process yields comparable results to coverage-based selection, highlighting that the second stage of Refine can easily adapted with more sophisticated components. We consider this a promising direction for future work, where such an selection mechanism could, for example, incorporate cost-sensitive criteria or realize a truly adaptive selection. For further details, we refer to our implementation.
Figure 11: Relative learning curves of Refine on Dopanim and CIFAR-100 (DINOv2) using coverage-based, FIR-based, and EER-based batch selection in the second stage.
Appendix D Additional Results and Experiments
Parallel Acquisition Times. To provide insights into the acquisition times of Refine and demonstrate how they can be reduced through parallelization, we report runtimes (min:sec) on CIFAR-100 (DINOv2) in Tab. 3. Sequential acquisition times remain on the same order as training times, and at later cycles the latter actually dominates. Since the selections of individual ensemble members are independent, parallelization is straightforward. We compute projected parallel runtimes using Amdahl’s Law, assuming a conservatively estimated 85% of the acquisition process is parallelizable, with a realistic scheduling overhead of 3% for 8 and 8% for 80 processes. With 8 processes, acquisition drops below training time, making it no longer the computational bottleneck. For reference, AutoAL [37], another computationally intensive ensemble AL method, requires 4:24 and 8:41 at cycles 10 and 20, respectively, on the same hardware. Unlike Refine, AutoAL cannot be parallelized as easily, as an auxiliary model is updated after all strategies have produced their selections, limiting scalability.
Table 3: Acquisition times (min:sec) of Refine on CIFAR-100 (DINOv2) at AL cycles 10 and 20.
[…]
10 07:50 09:58 02:38 01:44
20 14:34 10:31 02:47 01:50
Aggressiveness of Progressive Filtering. To provide insights into how aggressively progressive filtering reduces the size of the candidate pools ${\mathcal{C}}_{r}$, we report the ratio $\nicefrac{{|{\mathcal{C}}_{r}|}}{{|{\mathcal{U}}|}}$ across filtering rounds $r$ for several datasets in Tab. 4. For datasets with small batch sizes such as CIFAR-10 ($b=10$), filtering is highly aggressive: after five rounds, only approximately 5% of unlabeled instances remain in $\mathcal{C}_{R}$, meaning over 95% are discarded before final acquisition. For datasets with larger batch sizes such as CIFAR-100 and Tiny ImageNet, the reduction is notably less pronounced. This also connects to the batch size analysis in Fig. 10 (right), where larger batch sizes benefit Refine by considering a larger portion of the unlabeled pool.
Table 4: Ratio $\nicefrac{{|\mathcal{C}_{R}|}}{{|\mathcal{U}|}}$ across filtering rounds on several datasets.
Filtering Round 1 2 3 4 5
CIFAR-10 ($b=10$) 0.163 0.087 0.067 0.058 0.052
Dopanim ($b=50$) 0.453 0.293 0.232 0.199 0.179
CIFAR-100 ($b=100$) 0.689 0.415 0.323 0.277 0.247
Tiny ImageNet ($b=200$) 1.000 0.711 0.586 0.516 0.472
Robustness to Class Imbalance & Distribution Shift. To further investigate the robustness of Refine in more realistic settings, we conduct additional experiments considering both class imbalance and out-of-distribution data, i.e., data that strongly differs from the foundation model’s training distributions. For the class-imbalance setting, we run Refine (DINOv2) on long-tailed versions of Snacks and CIFAR-100 using exponential class decay with imbalance ratios $\text{IR}\in\{10,50\}$, denoting the ratio of the largest to the smallest class. For the out-of-distribution setting, we include BloodMNIST and DermaMNIST from the MedMNIST benchmark [39]. The results in Fig. 12 show that Refine maintains strong performance, demonstrating robustness across both settings.
@@ Appendix D Additional Results and Experiments
Figure 12: Absolute learning curves of ensemble AL methods on long-tailed and out-of-distribution datasets using DINOv2.
Fine-tuning Experiments. While the main experiments employ frozen features, reflecting current practice [13, 19, 20], Refine is agnostic to the downstream model. To verify this, we investigate the performance of ensemble AL methods when fine-tuning the last two transformer blocks of a DINOv2 pretrained transformer. This setting is more challenging, as training hyperparameters, i.e., learning rate and weight decay, had to be tuned on CIFAR-10 and transferred to the other datasets. Since the labeled pool grows in each AL cycle, these hyperparameters are highly likely to become suboptimal over the course of the AL process [25, 18]. The relative learning curves in Fig. 13 show that Refine maintains strong performance.
Figure 13: Relative learning curves of ensemble AL methods when fine-tuning the last two transformer blocks of DINOv2.
Appendix E Main Results: All Learning Curves
This section presents the complete set of learning curves used to derive Figs. 2 and 4, which were omitted from the main paper due to space limitations. For individual AL strategies, Fig. 14 displays the absolute learning curves (accuracy), while Fig. 15 shows the relative learning curves (accuracy relative to random sampling from ${\mathcal{U}}_{t}$). Similarly, for ensemble AL methods, Figs. 16 and 17 present the absolute and relative learning curves, respectively. These results provide a comprehensive overview supporting the conclusions drawn in the main text. Additionally, we include BoSS [21], an oracle strategy that approximates the optimal selection strategy by leveraging label information unavailable to standard AL strategies. By comparing against this oracle, we can assess how closely the AL strategies and Refine approach the theoretical optimum.
Figures 14 and 15 demonstrate that Refine outperforms all other individual AL strategies across nearly every dataset and backbone, including DINOv2, CLIP, and DINOv3. It excels in both low- and high-budget scenarios, consistently delivering high accuracies throughout the entire AL process. Furthermore, Refine proves to be robust, maintaining a strong lead across diverse datasets while remaining invariant to the choice of the backbone. Similarly, Figures 16 and 17 demonstrate that Refine outperforms all ensemble AL methods. These approaches generally struggle to effectively combine multiple strategies, often failing to exceed the performance of individual AL strategies. With Refine, we propose the first ensemble AL method capable of effectively combining multiple AL strategies into a unified framework that consistently surpasses individual baselines. This improvement over existing ensemble AL methods is statistically significant ($\alpha=0.01$), as confirmed by a Friedman test [10] on the AULC across 18 blocks (dataset $\times$ model) followed by post-hoc paired Wilcoxon signed-rank tests [38] with Holm correction [16].
When considering the optimal strategy approximated by the BoSS oracle, we observe that most strategies exhibit a considerable gap to this optimum. Notably, Refine closes this gap more effectively than other strategies, yet a significant gap to the oracle remains, underscoring the need for continued research into more effective AL selection strategies.
Figure 14: Complete set of absolute learning curves for individual AL strategies showing the accuracy for all datasets and backbone combinations. The optimal performance is approximated using BoSS [21].